# Evaluating Code-Mixing in LLMs Across 18 Languages

### v1

@@ Abstract
Code-mixing, the practice of switching between languages within a conversation, presents unique challenges for traditional natural language processing. Existing benchmarks, such as LinCE and GLUECoS, are limited by narrow language pairings and tasks, failing to adequately evaluate the code-mixing capabilities of large language models (LLMs). Despite the significance of code-mixing for multilingual users, research on LLMs in this context remains limited. Additionally, current methods for generating code-mixed data are underdeveloped. In this paper, we conduct a comprehensive evaluation of LLMs’ performance on code-mixed data across 18 languages from seven language families. We also propose a novel approach for generating synthetic code-mixed texts by combining word substitution with GPT-4 prompting. Our analysis reveals consistent underperformance of LLMs on code-mixed datasets involving multiple language families. We suggest that improvements in training data size, model scale, and few-shot learning could enhance their performance.
{CJK}
UTF8gkai
Evaluating Code-Mixing in LLMs Across 18 Languages
Yilun Yang NUC jeromeyluck@gmail.com                        Yekun Chaithanks: Corresponding author. Baidu chaiyekun@gmail.com
@@ 1 Introduction
Existing studies are difficult to compare directly because they focus on different downstream tasks and language pairs. To address this issue, LinCE (Aguilar et al., 2020) and GLUECoS (Khanuja et al., 2020b) introduced two code-switched benchmarks, but they cover a limited number of language pairs and traditional NLP tasks. LinCE addresses four language pairs and five traditional NLP tasks — language identification (LID), part-of-speech tagging (POS), named entity recognition (NER), sentiment analysis (SA), and machine translation (MT), while GLUECoS covers only two language pairs and six traditional NLP tasks — LID, POS, NER, SA, question answering (QA), and natural language inference (NLI). These traditional NLP tasks are inadequate for evaluating LLM performance effectively.
@@
Research on code-mixing in LLMs has been overlooked. Despite strong multilingual performance on various benchmarks, LLMs’ capabilities with code-mixing remain underexplored. Limited studies suggest that LLMs often perform worse than smaller, fine-tuned models on code-mixing tasks (Zhang et al., 2023), and multilingual users prefer chatbots that handle code-mixing well (Bawa et al., 2020). Thus, incorporating code-mixing into LLM evaluation is crucial.
Current techniques for synthesizing code-mixed data remain fundamentally limited in their effectiveness and scalability. The creation of high-quality code-mixed datasets for LLM evaluation continues to present significant challenges due to the inherent complexity of linguistic mixing phenomena. Prior approaches have primarily followed two methodological directions, each with substantial limitations. The first strand of research, grounded in Equivalence Constraint theory (POPLACK, 1980), has developed rule-based generation systems Bhat et al. (2016); Pratapa et al. (2018) that rely heavily on word alignment and syntactic parsing tools. However, these methods demonstrate critical weaknesses including their dependency on the accuracy of underlying NLP components, inability to capture authentic sociolinguistic mixing patterns, and poor scalability across diverse language pairs. More recent data-driven approaches employing neural models (Yang et al., 2020; Hsu et al., 2023) have achieved improved fluency but face distinct challenges related to their requirement for large existing code-mixed corpora, difficulties in maintaining grammaticality for low-resource languages, and tendencies to produce semantically inconsistent outputs. Notably, while large language models present a promising avenue for advancement, initial attempts to leverage their capabilities (Yong et al., 2023) have failed to fully exploit their instruction-following potential and contextual understanding.
In response to these issues, we introduce the CodeMixEval, , an analytical framework for diagnosing code-mixing capabilities in language models. Our study systematically examines performance across three key dimensions of LLM capabilities (knowledge reasoning, mathematical reasoning, and truthfulness) alongside five foundational NLP tasks (language identification, part-of-speech tagging, named entity recognition, sentiment analysis, and machine translation). This multilingual investigation spans 18 carefully selected languages representing seven distinct language families, including both high-resource and understudied language pairs. Our work focuses on revealing critical patterns in model behavior - particularly the relative strengths and limitations of current architectures when processing linguistically complex code-mixed inputs. The analysis enables detailed examination of cross-lingual interference effects, quantifies performance disparities between related and unrelated language pairs, and identifies promising directions for improving model robustness in real-world multilingual scenarios.
Our analytical framework extends beyond prior work by incorporating a more diverse set of language pairs (18 vs. 2-4 in existing studies) and introducing three novel LLM-specific evaluation dimensions (Appendix A).
@@
    We present CodeMixEval, the most extensive analytical framework to date for studying code-mixing in LLMs, systematically evaluating 8 task categories across 18 languages from 7 language families. Our methodology combines: (a) 22 newly synthesized datasets using our proposed pipeline, and (b) 30 carefully curated existing datasets (§3).
@@
    We develop a linguistically-informed synthesis pipeline that combines word substitution patterns with LLM prompting. Experimental results demonstrate this approach can generate contextually appropriate code-mixed text at scale, addressing key limitations of prior methods (§3.2).
@@
    We evaluate three families of LLMs on CodeMixEval, revealing consistent underperformance across all models on code-mixing datasets involving language pairs from different language families. However, enhancements in training data size, model scale, post-training, and few-shot learning can improve LLM performance on code-mixing datasets (§4).
@@ 2 Related Work
Early research employed linguistic rules and statistical methods (Li and Fung, 2012, 2014; Bhat et al., 2016; Rijhwani et al., 2017) for code-mixing modelling. Subsequently, research shifted towards neural network models like RNNs and LSTMs (Adel et al., 2013b, a; Wang et al., 2018; Winata et al., 2018), and more recently towards pre-trained language models such as mBERT and XLM-R (Winata et al., 2021; Malmasi et al., 2022; Pérez et al., 2022). These methodologies have been applied to various code-mixing-related downstream tasks, including language identification (Solorio et al., 2014; Molina et al., 2016), named entity recognition (Aguilar et al., 2018), part-of-speech tagging (Singh et al., 2018b; Soto and Hirschberg, 2018), sentiment analysis (Patra et al., 2018; Patwa et al., 2020), machine translation (Srivastava and Singh, 2020; Chen et al., 2022), natural language inference (Khanuja et al., 2020a), question answering (Chandu et al., 2018), and multilingual code generation Chai et al. (2023a); Peng et al. (2024). Benchmarks, such as GLUECoS Khanuja et al. (2020b) and LinCE Aguilar et al. (2020) primarily focus on traditional NLP tasks and are restricted to a limited number of languages. Recent research by Zhang et al. (2023) on the performance of multilingual LLMs in code-switching contexts indicates that, despite their strong capabilities across various monolingual tasks, they still yield inferior performance compared to fine-tuned smaller models.
@@
Early research synthesized code-mixed data based on linguistic rules. Following the EC theory, Bhat et al. (2016); Pratapa et al. (2018) utilized word alignment tools and syntactic parsers to enable the structural substitution and integration of lexical elements within aligned parse trees. Subsequently, researchers trained generative models to produce code-mixed data, such as a sequence-to-sequence model with a Pointer-Generator (Winata et al., 2019; Gupta et al., 2020), Generative Adversarial Networks (GAN) (Chang et al., 2019; Chai et al., 2021, 2023b), and Variational AutoEncoders (VAE) (Samanta et al., 2019b). An increasing number of works (Samanta et al., 2019a; Yang et al., 2020; Arora et al., 2023; Hsu et al., 2023) focused on extending pre-trained models for code-mixed data generation. Recently, Yong et al. (2023) examined the ability of LLMs to generate code-mixed text in Southeast Asian languages, finding that ChatGPT can produce fluent and natural code-mixed text for certain languages using a zero-shot prompting approach. Instead of using LLMs to directly generate code-mixed text, we revisit the EC theory and integrate its core principles into the prompt. Based on parallel corpora, we instruct the LLM to replace lexical elements between parallel sentences, thereby generating grammatically coherent code-mixed text.
3 CodeMixEval
@@
Table 1: Statistics of 18 languages from 7 families. Each language is assigned a unique code in this paper based on the ISO 639. The Pop. indicates the population in millions of speakers. The CC indicates ratios of languages in the CommomCrawl. The MSA and EA stand for Modern Standard Arabic and Egyptian Arabic.
@@ 3.1 Overview
To evaluate LLMs’ comprehension of multilingual code-mixed texts, we introduce CodeMixEval, a comprehensive evaluation framework comprising eight tasks across 18 languages. Table 1 details the speaker population and resource ratio on CommonCrawl [^1] https://commoncrawl.github.io/cc-crawl-statistics/plots/languages.html for each language, identified by their ISO 639 codes. The chosen languages exhibit diversity in language families, resource availability, and speaker populations. Motivated by Bang et al. (2023); Lai et al. (2023a, b), five languages (zh, es, fr, de, nl) are categorized as high-resource (CC >1%), three (ar, hi, bn) as mid-resource (0.1%-1%), and four (mr, ne, ta, ml) as low-resource (<0.01%).
@@ 4.1 Experiment Setup
Table 2: One-shot accuracy of selected models on CM-MMLU, CM-GSM8K and CM-TuthfulQA. Where 3.5-T indicates GPT-3.5 Turbo, and 4-T indicates GPT-4 Turbo. The en only stands for a dataset we randomly sample from the test set of the original dataset in English. To be compared with other code-mixed datasets, the en only datasets for CM-MMLU, CM-GSM8K, and CM-TuthfulQA contain 1000, 1133, and 817 English instances each. The Average represents the mean score of each model across various datasets (excluding en only dataset) from a given task. For each model family, the scores of the top-performing models are highlighted in bold.
Figure 2: One-shot accuracy versus language pairs for GPT models on LID, POS, NER, SA and MT.
@@ 4.2 Main Results
Larger models excel on CodeMixEval. In Table 2, GPT-4o achieves the highest scores across all language pairs in the CM-MMLU task, while GPT-4 Turbo attains the highest scores for each language pair in the CM-GSM8K and CM-TruthfulQA tasks. This suggests that GPT-4o excels in comprehensive knowledge reasoning, whereas GPT-4 Turbo is superior in mathematical reasoning and truthfulness. Additionally, within the LLaMA2, LLaMA3, and Mistral model families, the highest scores across all datasets consistently come from the largest models. Therefore, increasing model size enhances performance on multilingual code-mixed datasets.
GPT-3.5-Turbo-Instruct vs. GPT-3.5 Turbo. In Table 2, GPT-3.5 Turbo outperforms GPT-3.5-Turbo-Instruct by an average of 2.07 points in CM-MMLU, 14.54 points in CM-GSM8K, and 7.23 points in CM-TruthfulQA. Table 8 in the Appendix I shows GPT-3.5 Turbo scored higher on LID (+9.51%), POS (+1.68%), NER (+10.99%), and MT (+1.07%), but was 14.98 points lower on SA. This may be due to differing focuses during instruction tuning, with GPT-3.5 Turbo emphasizing conversational completion and GPT-3.5-Turbo-Instruct focusing on instruction completion, leading to different training corpora. Thus, GPT-3.5 Turbo excelled over GPT-3.5-Turbo-Instruct in all CodeMixEval tasks except SA.
LLaMA3-8B-Base vs. LLaMA3-8B-Instruct. In Table 2, LLaMA3-8B-Instruct performs comparably to LLaMA3-8B-Base on CM-MMLU and CM-TruthfulQA but outperforms it by 7.73 points on CM-GSM8K. This is likely due to the increased complexity of the mathematical reasoning required by CM-GSM8K. The improved performance on CM-GSM8K can be attributed to high-quality prompts during continued post-training stages, including supervised fine-tuning and alignment tuning, followed by the pre-training of LLaMA3.
LLaMA2 vs. LLaMA3. Table 2 shows that LLaMA3-8B outperforms LLaMA2-7B-Chat with average gains of 25.29, 57.29, and 23.43 points on CM-MMLU, CM-GSM8K, and CM-TruthfulQA, respectively. Additionally, LLaMA3-70B surpasses LLaMA2-70B with improvements of 25.75, 45.80, and 35.24 points on the same benchmarks. These enhancements may be due to the training dataset for LLaMA3 containing over 15T tokens, a size seven times larger than that used for LLaMA2.
Mistral 7B vs. Mixtral 8x7B. We also observe from Table 2 that Mixtral 8x7B outperforms Mistral 7B by 12.14, 19.73, and 8.98 points on CM-MMLU, CM-GSM8K, and CM-TruthfulQA, correspondingly. This improvement is likely due to the scaling of model parameters in Mixture of Experts (MoE) architecture and the substantial increase in multilingual training compute for Mixtral 8x7B.
@@ 4.4 $K$-shot Analysis
Few-shot learning minimally enhances mathematical reasoning. In Figure 4(b), the model’s $k$-shot accuracy on the CM-GSM8K task shows less variability compared to the other tasks, which we attribute to the higher complexity of CM-GSM8K relative to CM-MMLU and CM-TruthfulQA, posing greater challenges to a model’s few-shot learning. However, GPT-3.5 Turbo, GPT-3.5-Turbo-Instruct, and LLaMA3-8B-Base exhibit significant accuracy improvements from zero-shot to one-shot. This is because these models initially fail to follow the required response format in zero-shot prompts, causing incorrect answers, while one-shot improves these models’ output format adherence. Besides, Mixtral-7x22b demonstrates a decrease in accuracy as $k$ increases across all datasets, indicating its inadequate few-shot learning capability on CM-GSM8K. Overall, in the CM-GMS8K task, few-shot learning provides limited enhancement in mathematical reasoning for the GPT and LLaMA family and may negatively affect Mistral models.

### v2

# CodeMixBench: Evaluating Code-Mixing Capabilities of LLMs Across 18 Languages

@@ Abstract
Code-mixing, the practice of switching between languages within a conversation, poses unique challenges for traditional NLP. Existing benchmarks are limited by their narrow language pairs and tasks, failing to adequately assess large language models’ (LLMs) code-mixing abilities. Despite the recognized importance of code-mixing for multilingual users, research on LLMs in this context remains sparse. Additionally, current techniques for synthesizing code-mixed data are underdeveloped to generate code-mixing. In response, we introduce CodeMixBench, a comprehensive benchmark covering eight tasks, including three specific to LLMs and five traditional NLP tasks, and 18 languages across seven language families. We also propose a new method for generating large-scale synthetic code-mixed texts by combining word substitution with GPT-4 prompting. Our evaluation reveals consistent underperformance of LLMs on code-mixed datasets involving different language families. Enhancements in training data size, model scale, and few-shot learning could improve their performance. The code and dataset are available at https://github.com/Jeromeyluck/CodeMixBench.
CodeMixBench: Evaluating Code-Mixing Capabilities of LLMs
Across 18 Languages
Yilun Yang NUC jeromeyluck@gmail.com          Yekun Chaithanks: Corresponding author. ETH Zurich yechai@ethz.ch
@@ 1 Introduction
Existing studies are difficult to compare directly because they focus on different downstream tasks and language pairs. To address this issue, LinCE (Aguilar et al., 2020) and GLUECoS (Khanuja et al., 2020b) introduced two benchmarks, but they only cover a limited number of language pairs and traditional NLP tasks. LinCE addresses four language pairs and five traditional NLP tasks, including language identification (LID), part-of-speech tagging (POS), named entity recognition (NER), sentiment analysis (SA), and machine translation (MT), while GLUECoS covers only two language pairs and six traditional NLP tasks, i.e., LID, POS, NER, SA, question answering (QA), and natural language inference (NLI). These traditional NLP tasks are insufficient to evaluate LLM performance comprehensively.
@@
Despite strong multilingual performance on various benchmarks, LLMs’ capabilities with code-mixing remain underexplored. Limited studies suggest that LLMs often perform worse than smaller, fine-tuned models on code-mixing tasks (Zhang et al., 2023), and multilingual users prefer chatbots that handle code-mixing well (Bawa et al., 2020). Thus, incorporating code-mixing into LLM evaluation is crucial.
Creating new code-mixed datasets for LLMs involves using synthesis techniques. Some studies Bhat et al. (2016); Pratapa et al. (2018) focused on generating synthetic code-mixed data to solve the scarcity of code-mixed data, using methods based on the Equivalence Constraint theory (POPLACK, 1980), a linguistic theory that restricts the occurrences of code-mixing. However, the quality of these outputs heavily depends on the performance of word alignment and syntactic parsing tools. Recent efforts to generate code-mixed text using data-driven models still face challenges related to dataset size, quality, or linguistic diversity (Yang et al., 2020; Hsu et al., 2023). Also, initial attempts to use LLMs for generating code-mixed data did not fully leverage their instruction-following capabilities (Yong et al., 2023).
In response to these issues, we introduce the CodeMixBench, a code-mixing evaluation benchmark including eight tasks–three for evaluating LLMs (knowledge reasoning, mathematical reasoning, and truthfulness) and five for traditional NLP tasks (LID, POS, NER, SA, and MT). They span 18 languages from seven language families, covering high-resource, medium-resource, and low-resource languages. Our benchmark largely expands language pair and task coverage compared to LinCE and GLUECoS (Appendix A). We also propose a novel synthetic code-mixing approach using word substitution within GPT-4 prompting to generate large-scale code-mixed texts from parallel corpora.
@@
    We present CodeMixBench, the first comprehensive benchmark for evaluating the performance of LLMs on multilingual code-mixing. We have synthesized 22 datasets and, through extensive research, compiled 30 open-source code-mixed datasets to integrate into our benchmark. In total, the benchmark encompasses eight tasks and 18 languages from seven language families (§3).
@@
    We propose a novel pipeline for large-scale synthesis of multilingual code-mixing data, integrating word substitution with LLM prompts for the first time. The synthetic results validate the efficiency of our approach in generating substantial multilingual code-mixed data. (§3.2).
@@
    We evaluate three families of LLMs on CodeMixBench, revealing consistent underperformance across all models on code-mixing datasets involving language pairs from different language families. However, enhancements in training data size, model scale, post-training, and few-shot learning can improve LLM performance on code-mixing datasets (§4).
@@ 2 Related Work
Early research employed linguistic rules and statistical methods (Li and Fung, 2012, 2014; Bhat et al., 2016; Rijhwani et al., 2017) for code-mixing modeling. Subsequently, research shifted towards neural network models like RNNs and LSTMs (Adel et al., 2013b, a; Wang et al., 2018; Winata et al., 2018), and more recently towards pre-trained language models such as mBERT and XLM-R (Winata et al., 2021; Malmasi et al., 2022; Pérez et al., 2022). These methodologies have been applied to various code-mixing-related downstream tasks, including language identification (Solorio et al., 2014; Molina et al., 2016), named entity recognition (Aguilar et al., 2018), part-of-speech tagging (Singh et al., 2018b; Soto and Hirschberg, 2018), sentiment analysis (Patra et al., 2018; Patwa et al., 2020), machine translation (Srivastava and Singh, 2020; Chen et al., 2022), natural language inference (Khanuja et al., 2020a), question answering (Chandu et al., 2018), and multilingual code generation Chai et al. (2023a); Peng et al. (2024). Benchmarks, such as GLUECoS Khanuja et al. (2020b) and LinCE Aguilar et al. (2020) primarily focus on traditional NLP tasks and are restricted to a limited number of languages. Recent research by Zhang et al. (2023) on the performance of multilingual LLMs in code-switching contexts indicates that, despite their strong capabilities across various monolingual tasks, they still yield inferior performance compared to fine-tuned smaller models.
@@
Early research synthesized code-mixed data based on linguistic rules. Following the EC theory, Bhat et al. (2016); Pratapa et al. (2018) utilized word alignment tools and syntactic parsers to enable the structural substitution and integration of lexical elements within aligned parse trees. Subsequently, researchers trained generative models to produce code-mixed data, such as a sequence-to-sequence model with a Pointer-Generator (Winata et al., 2019; Gupta et al., 2020), Generative Adversarial Networks (Chang et al., 2019; Chai et al., 2021, 2023b), and Variational AutoEncoders (Samanta et al., 2019b). An increasing number of works (Samanta et al., 2019a; Yang et al., 2020; Arora et al., 2023; Hsu et al., 2023) focused on extending pre-trained models for code-mixed data generation. Yong et al. (2023) examined the ability of LLMs to generate code-mixed text in Southeast Asian languages. Instead of using LLMs to directly generate code-mixed text, we revisit the EC theory and integrate its core principles into the prompt. Based on parallel corpora, we instruct the LLM to replace lexical elements between parallel sentences, thereby generating grammatically coherent code-mixed text.
3 CodeMixBench
@@
Table 1: Statistics of 18 languages from 7 families. Each language is assigned a unique code in this paper based on the ISO 639. The Pop. indicates the population in millions of speakers. The $C$ indicates ratios of languages in the CommomCrawl. The MSA and EA stand for Modern Standard Arabic and Egyptian Arabic.
@@ 3.1 Overview
To evaluate LLMs’ comprehension of multilingual code-mixed texts, we introduce CodeMixBench, a benchmark comprising eight tasks across 18 languages. Table 1 details the speaker population and resource ratio on CommonCrawl [^1] https://commoncrawl.github.io/cc-crawl-statistics/plots/languages.html for each language, identified by their ISO 639 codes. The chosen languages exhibit diversity in language families, resource availability, and speaker populations. Motivated by Bang et al. (2023); Lai et al. (2023a, b), five languages (zh, es, fr, de, nl) are categorized as high-resource (CC >1%), three (ar, hi, bn) as mid-resource (0.1%-1%), and four (mr, ne, ta, ml) as low-resource (<0.01%).
@@ 4.1 Experiment Setup
Table 2: One-shot accuracy of selected models on CM-MMLU, CM-GSM8K and CM-TruthfulQA. Where 3.5-T indicates GPT-3.5 Turbo, and 4-T indicates GPT-4 Turbo. The en only stands for a dataset we randomly sample from the test set of the original dataset in English. To be compared with other code-mixed datasets, the en only datasets for CM-MMLU, CM-GSM8K, and CM-TuthfulQA contain 1000, 1133, and 817 English instances each. The Average represents the mean score of each model across various datasets (excluding en only dataset) from a given task. For each model family, the scores of the top-performing models are highlighted in bold.
Figure 2: One-shot accuracy versus language pairs for GPT models on LID, POS, SA, NER and MT.
@@ 4.2 Main Results
Larger models excel on CodeMixBench
In Table 2, GPT-4o achieves the highest scores across all language pairs in the CM-MMLU task, while GPT-4 Turbo attains the highest scores for each language pair in the CM-GSM8K and CM-TruthfulQA tasks. This suggests that GPT-4o excels in comprehensive knowledge reasoning, whereas GPT-4 Turbo is superior in mathematical reasoning and truthfulness. Additionally, within the LLaMA2, LLaMA3, and Mistral model families, the highest scores across all datasets consistently come from the largest models. Therefore, increasing model size enhances performance on multilingual code-mixed datasets.
GPT-3.5-Turbo-Instruct vs. GPT-3.5 Turbo
In Table 2, GPT-3.5 Turbo outperforms GPT-3.5-Turbo-Instruct by an average of 2.07 points in CM-MMLU, 14.54 points in CM-GSM8K, and 7.23 points in CM-TruthfulQA. Table 8 in the Appendix I shows GPT-3.5 Turbo scored higher on LID (+9.51%), POS (+1.68%), NER (+10.99%), and MT (+1.07%), but was 14.98 points lower on SA. This may be due to differing focuses during instruction tuning, with GPT-3.5 Turbo emphasizing conversational completion and GPT-3.5-Turbo-Instruct focusing on instruction completion, leading to different training corpora. Thus, GPT-3.5 Turbo excelled over GPT-3.5-Turbo-Instruct in all CodeMixBench tasks except SA.
LLaMA3-8B-Base vs. LLaMA3-8B-Instruct
In Table 2, LLaMA3-8B-Instruct performs comparably to LLaMA3-8B-Base on CM-MMLU and CM-TruthfulQA but outperforms it by 7.73 points on CM-GSM8K. This is likely due to the increased complexity of the mathematical reasoning required by CM-GSM8K. The improved performance on CM-GSM8K can be attributed to high-quality prompts during continued post-training stages, including supervised fine-tuning and alignment tuning, followed by the pre-training of LLaMA3.
LLaMA2 vs. LLaMA3
Table 2 shows that LLaMA3-8B outperforms LLaMA2-7B-Chat with average gains of 25.29, 57.29, and 23.43 points on CM-MMLU, CM-GSM8K, and CM-TruthfulQA, respectively. Additionally, LLaMA3-70B surpasses LLaMA2-70B with improvements of 25.75, 45.80, and 35.24 points on the same benchmarks. These enhancements may be due to the training dataset for LLaMA3 containing over 15T tokens, a size seven times larger than that used for LLaMA2.
Mistral 7B vs. Mixtral 8x7B
We also observe from Table 2 that Mixtral 8x7B outperforms Mistral 7B by 12.14, 19.73, and 8.98 points on CM-MMLU, CM-GSM8K, and CM-TruthfulQA, correspondingly. This improvement is likely due to the scaling of model parameters in Mixture of Experts (MoE) architecture and the substantial increase in multilingual training compute for Mixtral 8x7B.
@@ 4.4 $K$-shot Analysis
Few-shot learning minimally enhances mathematical reasoning. In Figure 4(b), the model’s $k$-shot accuracy on the CM-GSM8K task shows less variability compared to the other tasks, which we attribute to the higher complexity of CM-GSM8K relative to CM-MMLU and CM-TruthfulQA, posing greater challenges to a model’s few-shot learning. However, GPT-3.5 Turbo, GPT-3.5-Turbo-Instruct, and LLaMA3-8B-Base exhibit significant accuracy improvements from zero-shot to one-shot. This is because these models initially fail to follow the required response format in zero-shot prompts, causing incorrect answers, while one-shot improves these models’ output format adherence. Besides, Mixtral-7x22b demonstrates a decrease in accuracy as $k$ increases across all datasets, indicating its inadequate few-shot learning capability on CM-GSM8K. Overall, in the CM-GSM8K task, few-shot learning provides limited enhancement in mathematical reasoning for the GPT and LLaMA family and may negatively affect Mistral models.